OpenStudy (elusive):

how would you write 1/ 1/2 ?

OpenStudy (mariamikayla):

1/1/2= 1 : 1/2 = 1 * 2/1= 2

OpenStudy (elusive):

\[\frac{ 1 }{ \sqrt{3}/2 }\] how about this?

OpenStudy (elusive):

@Mariamikayla It doesn't make sense.. Shouldn't that be 0.5? Half of 1..

OpenStudy (mariamikayla):

how your first fraction is written? as 1 on 1/2 or 1/1 on 2?

OpenStudy (elusive):

\[\frac{ 1 }{ 1/2 }\]

OpenStudy (anthonyym):

It is 1 divided by 1/2, so how many times does 1/2 go into 1? Twice.

OpenStudy (mariamikayla):

1/ root 3/2= 1: root 3/2= 1* 2/root 3= 2/root 3 we rationalize and we obtain ---> 2 root 3/ 3

OpenStudy (elusive):

\[\frac{ 1 }{ \sqrt{2}/2 } = \frac{ 2 }{ \sqrt{2} }\] so is this also correct?

OpenStudy (mariamikayla):

yes